# Pulsar timing array sensitivity to anisotropies in the gravitational wave background

### v1

@@ I Introduction
Recent results from pulsar timing arrays (PTAs) around the world have established a first indication of gravitational waves in the nanohertz range NANOGrav:2023gor ; Antoniadis:2023ott ; Reardon:2023gzh ; Xu:2023wog . The observed signal is consistent with a gravitational wave background (GWB), showing in particular evidence of an angular correlation following the Hellings-Downs (HD) function Hellings:1983fr , considered a smoking gun signal for gravitational wave (GW) detection at PTAs. The origin of this signal is less clear. While the most likely explanation remains a GWB emitted by supermassive black hole binaries (SMBHBs), a range of scenarios giving rise to a cosmological GWB sourced in the early Universe are currently also compatible with the data, see, e.g., Madge:2023cak ; NANOGrav:2023hvm ; Antoniadis:2023xlr ; Figueroa:2023zhu ; Ellis:2023oxs . A key to distinguishing these two options lies in the search for anisotropies in the signal: while most cosmological backgrounds are expected to be highly isotropic, the GWB from SMBHBs is expected to feature anisotropies at the level of $1-20\%$ at low multipoles Mingarelli:2013dsa ; Taylor:2013esa ; Mingarelli:2017fbe ; NANOGrav:2023tcn . This raises the question of how well upcoming PTA data sets are expected to measure this quantity.
@@
The remainder of this paper is organized as follows. Section II introduces the theoretical framework describing the impact of anisotropies in the GWB on all the auto- and cross-correlations of pulsar timing delays. Based on this, Sec. V.4 introduces the FIM formalism which can be used to estimate and forecast PTA sensitivities to anisotropic GW signals. We derive the key properties of our main results semi-analytically in the strong signal limit in Sec. IV before presenting our main numerical findings as well as their interpretation in Sec. V. We conclude in Sec. VI. Three appendices are dedicated to the derivation of the covariance matrix, the explicit expressions and properties of the generalized overlap reduction functions, and further comparisons with the NG15 results.
@@ II.1 Stochastic Gravitational Wave Background
and normalized such that $\int{\rm d}^{2}\Omega_{\hat{k}}Y_{\ell m}Y_{\ell^{\prime}m^{\prime}}=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$. Additionally, we impose that
@@
With $Y_{00}=1/\sqrt{4\pi}$ and $\int{\rm d}^{2}\Omega_{\hat{k}}Y_{\ell m}=0$ for $\ell\geq 1$, the normalization condition $\int{\rm d}^{2}\Omega_{\hat{k}}P(\hat{k})=1$ imposes $c_{00}=1/\sqrt{4\pi}$. The PSD of an isotropic GWB is thus given directly by $S_{h}(f)$.
@@ II.2 Pulsar Timing Array Response
PTAs measure the timing residual $\Delta t_{I}$ induced by GWs along the line-of-sight between a pulsar $I$ (located in direction $\hat{p}_{I}$ at distance $D$) and Earth (located here at position $\vec{0}$),
$\Delta t_{I}=\frac{1}{2}\frac{\hat{p}_{I}^{a}\hat{p}_{I}^{b}}{(1+\hat{p}_{I}\cdot\hat{k})}\int_{0}^{t}{\rm dt^{\prime}}\left[h_{ab}(t^{\prime},\vec{0})-h_{ab}(t^{\prime}-D,\hat{p}_{I}D)\right].$ (7)
In practice, when extracting the timing residuals by fitting a pulsar timing model we need to additionally include a transmission function, see below. Expanding GWs as in Eq. (1) yields
$\Delta t_{I}(t)=\int_{-\infty}^{\infty}{\rm d}f\>\int{\rm d}^{2}\Omega_{\hat{k}}\>\sum_{P}R_{I}^{P}(f,\hat{k})\,\tilde{h}_{P}(f;\hat{k})\left[e^{2\pi ift}-1\right]\,,$ (8)
@@
$R_{I}^{P}(f,\hat{k})\equiv\frac{F^{P}_{\hat{p}_{I}}(\hat{k})}{i2\pi f}\left(1-e^{-i2\pi fD(1+\hat{k}\cdot\hat{p}_{I})}\right)\qquad\text{with}\qquad F^{P}_{\hat{p}_{I}}(\hat{k})\equiv\frac{1}{2}\frac{\hat{p}^{a}_{I}\hat{p}_{I}^{b}}{1+\hat{p}_{I}\cdot{\hat{k}}}e^{P}_{ab}({\hat{k}})\,.$ (9)
Here we have introduced the response function $R_{I}^{P}(f,\hat{k})$, capturing the impact of a GW with frequency $f$, polarization $P$ and propagating in the $\hat{k}$ direction on the signal from the $I$th pulsar. We note that the unknown GW phase at $t=0$ in Eq. (8) together with the convention $\Delta t(t=0)=0$ leads, in general, to an oscillation of $\Delta t$ around a non-zero value. We comment on this point below. The correlation function of timing residuals from pulsars $I$ and $J$ observed at times $t_{i}$ and $t_{j}$ is then
$C_{t,h,IJ,ij}$ $=\langle\Delta t_{I}(t_{i})\Delta t_{J}(t_{j})\rangle=\int_{-\infty}^{\infty}{\rm d}f\>\Gamma_{IJ}(f)\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(e^{2\pi ift_{i}}-1\right)\left(e^{2\pi ift_{j}}-1\right)\,,$ (10)
@@
To have the correlation coefficient equal to unity if the two pulsars are identical, we have collected an overall factor $3/2$ Romano:2023zhb in the definition of $\gamma_{IJ}$. [^1] Notice also a factor $4$ difference compared to the notation used in Ali-Haimoud:2020ozu for the pairwise response function.
To simplify this expression we notice that for PTAs, the minimum frequency is limited by the observation time $f\sim 1/T_{\rm obs}\sim\text{nHz}\sim 1/\text{pc}$. Moreover, the typical distance to the pulsars is $D\sim~{}\text{kpc}$. Therefore, $fD\gg 1$, and the rapidly oscillating pieces in Eq. (12) are negligible when integrated over the directions $\hat{k}$ (see also Mingarelli:2013dsa ). As a consequence, we can drop the terms proportional to the rapidly oscillating ‘pulsar terms’ in $\kappa_{IJ}$. Under these assumptions we have,
@@
On the other hand, in the more general case, anisotropies in the GW background, encoded in $P(\hat{k})$, are convoluted with the angular response function $\gamma_{IJ}$ in Eq. (13) and imprinted onto the correlation pattern of the time delays. Expanding also the instrument response in terms of spherical harmonics with coefficients $\Gamma_{IJ,\ell m}$,
@@ II.3 Covariance Matrix in the Frequency Domain
$\simeq T_{\text{obs}}^{2}\int_{-\infty}^{\infty}{\rm d}f\>\Gamma_{IJ}(f)\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(\mathrm{sinc}(\pi T_{\text{obs}}(f-f_{k}))-\delta_{k0}\right)\left(\mathrm{sinc}(\pi T_{\text{obs}}(f-f_{l}))-\delta_{l0}\right)\,.$ (20)
The terms $\delta_{k0}$ and $\delta_{l0}$ appear from the integration at $t^{\prime}=0$ in Eq. (7) and are removed from our Fisher analysis, thus removing the boundary term. For $k,l\neq 0$, we can evaluate the integral over $f$ above explicitly. At this point, we also switch to using only positive frequencies, $f_{k},f_{l}>0$. Then, if $\Gamma_{IJ}(f)S_{h}(f)/f^{2}$ is approximately constant for frequencies $f-f_{k}\lesssim 1/T_{\text{obs}}$ and $f-f_{l}\lesssim 1/T_{\text{obs}}$, one may approximate $\mathrm{sinc}(\pi T_{\text{obs}}[f-f_{k}])\mathrm{sinc}(\pi T_{\text{obs}}[f-f_{l}])\simeq 2\delta_{kl}\delta(f-f_{k})/T_{\text{obs}}$ leading to
$C_{h,IJ,kl}$ $\simeq\Gamma_{IJ}(f_{k})\frac{S_{h}(f_{k})}{12\pi^{2}f_{k}^{2}}\delta(f_{k}-f_{l})\,.$ (21)
For a discussion of the accurateness of these approximations see appendix A.
@@ III Fisher forecast of sensitivities
$-\ln\mathcal{L}(\tilde{d}|{\theta})\propto\sum_{k,IJ}\ln\left[C_{IJ}(f_{k},{\theta})\right]+\tilde{d}_{I}^{k}C_{IJ}^{-1}(f_{k},{\theta})\tilde{d}_{J}^{k*},$ (26)
@@
$F_{\alpha\beta}=\sum_{f_{k}}C^{-1}_{IJ}C^{-1}_{KL}\frac{\partial C_{h,JK}}{\partial\theta_{\alpha}}\frac{\partial C_{h,LI}}{\partial\theta_{\beta}}\,.$ (27)
@@ V.4 Cosmic Variance
PTA experiments operate in a regime where neither the time average over observation time nor the volume average over the effective detector volume, set by the wavelength of the GWs, allows sampling many realizations of a GWB. Consequently, measurements may be sizeably affected by cosmic variance. This notably affects the measurement of the angular cross-correlation (HD curve), see Refs. Allen:2022dzg ; Allen:2022ksj ; Allen:2024uqs . In this subsection, we clarify the role of cosmic variance in measuring anisotropies. For simplicity, we will focus on the strong signal limit.
@@ V.5 Hellings-Downs correlation
It is well known that for an anisotropic GWB, the observed angular correlation function scatters around the HD function (see e.g. Taylor:2013esa ). After averaging over many pulsar pairs, the HD function is recovered as the mean value [^9] This implies taking an ensemble average, whereas any actual measurement will be limited by cosmic variance. of the angular correlation function Hotinli:2019tpc , however with an increased variance reflecting the presence of anisotropies Agarwal:2024hlj ; Allen:2024bnk ; Grimm:2024lfj ; Allen:2024mtn . This implies on the one hand, that the measurement of the HD correlation (averaged over many pulsars) is largely insensitive to the presence of anisotropies. On the other hand, this implies that the HD function is a robust criterion for a GW signal, both for isotropic GWBs, isolated sources, and anisotropic signals arising from multiple point sources Cornish:2013aba ; Cornish:2015ikx ; Taylor:2020zpk ; Becsy:2022pnr ; Allen:2022dzg .
@@ Acknowledgements.
We thank Nihan Pol for the useful discussion and clarifications on the NANOGrav search for anisotropies and Fabrizio Rompineve for collaboration on the initial stages of the project. We thank Bruce Allen for valuable discussions on the HD cosmic variance. Moreover, we thank Stanislav Babak, and Carlo Contaldi for the insightful discussions. We thank Ben Farr and Andrea Mitridate for comments on the draft. MP thanks Rutger van Haasteren for the useful discussion on PTA data analysis. MP acknowledges the hospitality of Imperial College London, which provided office space during some parts of this project.
@@ Appendix A Approximations in the derivation of the covariance matrix in the frequency domain
In the following, we assume that there are $N$ observations at $t_{j}=-T_{\text{obs}}/2+jT_{\text{obs}}/N$, $j=0,\dots,N-1$, of all $\Delta t_{I}$ with pulsar index $I$ from 0 to $N_{p}$. As discussed in Sec. II.2, the complete covariance matrix is then constructed from
$C_{t,h,IJ,ij}$ $=\langle\Delta t_{I}(t_{i})\Delta t_{J}(t_{j})\rangle=\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(e^{2\pi ift_{i}}-1\right)\left(e^{2\pi ift_{j}}-1\right)\,,$
where $\chi_{IJ}$ is the HD correlation reported in Eq. (15). In the case of anisotropies in the GWB, this correlation must be replaced by the overlap reduction function $\Gamma_{IJ}$ given in Eq. (11). Defining the vector of all timing residuals
@@
$=\frac{T_{\text{obs}}^{2}}{N^{2}}\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(\frac{e^{-i\pi T_{\text{obs}}(f-f_{k})}-e^{i\pi T_{\text{obs}}(f-f_{k})}}{1-e^{i2\pi T_{\text{obs}}(f-f_{k})/N}}-N\delta_{k0}\right)\left(\frac{e^{i\pi T_{\text{obs}}(f-f_{l})}-e^{-i\pi T_{\text{obs}}(f-f_{l})}}{1-e^{-i2\pi T_{\text{obs}}(f-f_{l})/N}}-N\delta_{l0}\right)\,.$ (41)
In the low-frequency regime with large $N$, i.e., for $T_{\text{obs}}(f-f_{k})/N\ll 1$ and $T_{\text{obs}}(f-f_{l})/N\ll 1$, we can approximate
$C_{h,IJ,kl}$ $\simeq T_{\text{obs}}^{2}\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(\mathrm{sinc}(\pi T_{\text{obs}}(f-f_{k}))-\delta_{k0}\right)\left(\mathrm{sinc}(\pi T_{\text{obs}}(f-f_{l}))-\delta_{l0}\right)\,.$ (42)
The factors involving $\delta_{k0}$ and $\delta_{l0}$ appear due to the unknown GW phase at $t=0$, leading to the “-1” term in Eq. (8). Clearly, this term only contributes to the zero-frequency modes in the discrete Fourier transform. Restricting to positive frequencies, $k,l>0$ and if $S_{h}(f)/f^{2}$ is approximately constant for frequencies $f-f_{k}\lesssim 1/T_{\text{obs}}$ and $f-f_{l}\lesssim 1/T_{\text{obs}}$ one may approximate $\mathrm{sinc}(\pi T_{\text{obs}}[f-f_{k}])\mathrm{sinc}(\pi T_{\text{obs}}[f-f_{l}])\simeq 2\delta_{kl}\delta(f-f_{k})/T_{\text{obs}}$ leading to the expression used in the main text,

### v2

@@ I Introduction
Recent results from pulsar timing arrays (PTAs) around the world have established a first indication of gravitational waves in the nanohertz range NANOGrav:2023gor ; Antoniadis:2023ott ; Reardon:2023gzh ; Xu:2023wog . The observed signal is consistent with a gravitational wave background (GWB), showing in particular evidence of an angular correlation following the Hellings-Downs (HD) function Hellings:1983fr , considered a smoking gun signal for gravitational wave (GW) detection at PTAs. The origin of this signal is less clear. While the most likely explanation remains a GWB emitted by supermassive black hole binaries (SMBHBs), a range of scenarios giving rise to a cosmological GWB sourced in the early Universe are currently also compatible with the data, see, e.g., Madge:2023cak ; NANOGrav:2023hvm ; Antoniadis:2023xlr ; Figueroa:2023zhu ; Ellis:2023oxs . A key to distinguishing these two options lies in the search for anisotropies in the signal: while most cosmological backgrounds are expected to be highly isotropic, the GWB from SMBHBs is expected to feature anisotropies at the level of $1-20\%$ at low multipoles Mingarelli:2013dsa ; Taylor:2013esa ; Mingarelli:2017fbe ; NANOGrav:2023tcn ; Sah:2024oyg ; Sah:2024etc . This raises the question of how well upcoming PTA data sets are expected to measure this quantity.
@@
The remainder of this paper is organized as follows. Section II introduces the theoretical framework describing the impact of anisotropies in the GWB on all the auto- and cross-correlations of pulsar timing delays. Based on this, Sec. III introduces the FIM formalism, which can be used to estimate and forecast PTA sensitivities to anisotropic GW signals. We derive the key properties of our main results semi-analytically in the strong signal limit in Sec. IV before presenting our main numerical findings as well as their interpretation in Sec. V. We conclude in Sec. VI. Three appendices are dedicated to the derivation of the covariance matrix, the explicit expressions and properties of the generalized overlap reduction functions, and further comparisons with the NG15 results.
@@ II.1 Stochastic Gravitational Wave Background
and normalized such that $\int{\rm d}^{2}\Omega_{\hat{k}}Y_{\ell m}Y_{\ell^{\prime}m^{\prime}}=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}$. Notice that the normalization condition $\int{\rm d}^{2}\Omega_{\hat{k}}P(\hat{k})=1$ reads
@@
Since $Y_{00}=1/\sqrt{4\pi}$ and $\int{\rm d}^{2}\Omega_{\hat{k}}Y_{\ell m}=0$ for $\ell\geq 1$, this imposes $c_{00}=1/\sqrt{4\pi}$. The PSD of an isotropic GWB is thus given directly by $S_{h}(f)$.
@@ II.2 Pulsar Timing Array Response
PTAs measure the timing residual $\Delta t_{I}$ induced by GWs along the line-of-sight between a pulsar $I$ (located in direction $\hat{p}_{I}$ at distance $D_{I}$) and Earth (located here at position $\vec{0}$). The time shift induced on a photon reaching Earth at time $t$ then reads
$\Delta t_{I}=\frac{1}{2}\frac{\hat{p}_{I}^{a}\hat{p}_{I}^{b}}{(1+\hat{p}_{I}\cdot\hat{k})}\int_{0}^{D_{I}}{\rm d}s\,h_{ab}\left(t(s),\vec{x}(s)\right)\,,$ (7)
with $s$ denoting the affine parameter parameterising the geodesic connecting and $t(s)=t-(D_{I}-s)$, $\vec{x}=(D_{I}-s)\hat{p}_{I}$. In practice, when extracting the timing residuals by fitting a pulsar timing model we need to additionally include a transmission function Hazboun:2019vhv , see discussion around Eq. (22). Expanding GWs as in Eq. (1) yields
$\Delta t_{I}(t)=\int_{-\infty}^{\infty}{\rm d}f\>\int{\rm d}^{2}\Omega_{\hat{k}}\>\sum_{P}R_{I}^{P}(f,\hat{k})\,\tilde{h}_{P}(f;\hat{k})\,e^{2\pi ift}\,,$ (8)
@@
$R_{I}^{P}(f,\hat{k})\equiv\frac{F^{P}_{\hat{p}_{I}}(\hat{k})}{i2\pi f}\left(1-e^{-i2\pi fD_{I}(1+\hat{k}\cdot\hat{p}_{I})}\right)\qquad\text{with}\qquad F^{P}_{\hat{p}_{I}}(\hat{k})\equiv\frac{1}{2}\frac{\hat{p}^{a}_{I}\hat{p}_{I}^{b}}{1+\hat{p}_{I}\cdot{\hat{k}}}e^{P}_{ab}({\hat{k}})\,.$ (9)
Here we have introduced the response function $R_{I}^{P}(f,\hat{k})$, capturing the impact of a GW with frequency $f$, polarization $P$ and propagating in the $\hat{k}$ direction on the signal from the $I$th pulsar. The correlation function of timing residuals from pulsars $I$ and $J$ observed at times $t_{i}$ and $t_{j}$ is then
$C_{t,h,IJ,ij}$ $=\langle\Delta t_{I}(t_{i})\Delta t_{J}(t_{j})\rangle=\int_{-\infty}^{\infty}{\rm d}f\>\Gamma_{IJ}(f)\frac{S_{h}(f)}{24\pi^{2}f^{2}}e^{2\pi if(t_{i}-t_{j})}\,,$ (10)
@@
To have the overlap reduction function equal to unity if the two pulsars are identical, we have collected an overall factor $3/2$ Romano:2023zhb in the definition of $\gamma_{IJ}$. [^1] Notice also a factor $4$ difference compared to the notation used in Ali-Haimoud:2020ozu for the pairwise response function. In the discussion above, we have implicitly assumed the timing model to be unbiased, i.e. calibrated to Minkowski space time. In practice, this may be challenging to accomplish at low frequencies. We demonstrate in App. A that this does not significantly impact the analysis in the PTA frequency band.
To simplify Eq. (10), we notice that for PTAs, the minimum frequency is limited by the observation time $f\sim 1/T_{\rm obs}\sim\text{nHz}\sim 0.1/\text{pc}$. Moreover, the typical distance to the pulsars is $D_{I}\sim~{}\text{kpc}$. Therefore, $fD_{I}\gg 1$, and the rapidly oscillating pieces in Eq. (12) are negligible when integrated over the directions $\hat{k}$ (see also Mingarelli:2013dsa ). As a consequence, we can drop the terms proportional to the rapidly oscillating ‘pulsar terms’ in $\kappa_{IJ}$. Under these assumptions we have,
@@
On the other hand, in the more general case, anisotropies in the GW background, encoded in $P(\hat{k})$, are convolved with the angular response function $\gamma_{IJ}$ in Eq. (13) and imprinted onto the correlation pattern of the time delays. Expanding also the instrument response in terms of spherical harmonics with coefficients $\Gamma_{IJ,\ell m}$,
@@ II.3 Covariance Matrix in the Frequency Domain
$\simeq T_{\text{obs}}^{2}\int_{-\infty}^{\infty}{\rm d}f\>\Gamma_{IJ}\frac{S_{h}(f)}{24\pi^{2}f^{2}}\,\mathrm{sinc}\left[\pi T_{\text{obs}}(f-f_{k})\right]\,\mathrm{sinc}\left[\pi T_{\text{obs}}(f-f_{l})\right]\,.$ (20)
If $S_{h}(f)/f^{2}$ is approximately constant, this integration can be performed analytically. At this point we can also switch to using positive frequencies $f_{k}=k/T_{\text{obs}}$ only and obtain,
$C_{h,IJ,kl}$ $\simeq\Gamma_{IJ}(f_{k})\frac{S_{h}(f_{k})}{12\pi^{2}f_{k}^{2}}\,T_{\text{obs}}\,\mathrm{sinc}[\pi T_{\text{obs}}(f_{k}-f_{l})]\simeq\Gamma_{IJ}(f_{k})\frac{S_{h}(f_{k})}{12\pi^{2}f_{k}^{2}}\delta(f_{k}-f_{l})\,.$ (21)
For a discussion of the accurateness of these approximations, in particular regarding possible correlations across different frequencies, see appendix A.
@@ III Fisher forecast of sensitivities
$-\ln\mathcal{L}(\tilde{d}|{\theta})={\rm const.}+\sum_{k,IJ}\ln\left[C_{IJ}(f_{k},{\theta})\right]+\tilde{d}_{I}^{k}C_{IJ}^{-1}(f_{k},{\theta})\tilde{d}_{J}^{k*}\;,$ (26)
@@
$F_{\alpha\beta}=\sum_{f_{k},IJKL}C^{-1}_{IJ}C^{-1}_{KL}\frac{\partial C_{h,JK}}{\partial\theta_{\alpha}}\frac{\partial C_{h,LI}}{\partial\theta_{\beta}}\,.$ (27)
@@ V.3 Forecasts for future PTA configurations
V.3.1 Kinematic dipole
One inevitable deviation from isotropy is caused by the motion of the solar system barycenter (our local system) relative to the cosmological rest frame. Observation of the cosmic microwave background radiation indicate that our local system is traveling at a velocity of $\beta_{\rm proper}=v/c=1.2\cdot 10^{-3}$ towards the direction specified by galactic coordinates $(l=264^{\circ},b=48^{\circ})$. Assuming to have a similar velocity compared to the GWB rest frame, the anisotropies resulting from our proper motion can be described by
$\sum_{\ell m}c_{\ell m}Y_{\ell m}(\theta,\phi)=1+\beta_{\rm proper}(\cos\theta\cos 97^{\circ}+\sin\theta\sin 97^{\circ}\cos\phi).$ (34)
where the angles indicate spherical coordinates in the earth frame. One thus obtains the following multipole moments
$c_{00}$ $=1/\sqrt{4\pi};\qquad c_{1-1}=-c_{11}=\beta_{\rm proper}\sqrt{1\over{6}}\sin 97^{\circ};\qquad c_{10}=\beta_{\rm proper}\sqrt{1\over 3}\cos 97^{\circ};\qquad c_{\ell m}=0\quad{\rm for\ }\ell>1.$ (35)
In the strong signal limit described in Sec. IV, we find that the $1\sigma$ uncertainty on $\beta_{\rm proper}$ scale as
$\Delta\beta_{\rm proper}\simeq 0.8\times 1/\sqrt{N_{p}\,N_{f}}\qquad\quad(\text{strong signal limit}).$ (36)
This means that, in order to reach detection capability at around 90% C.L., $N_{f}=10^{2}$ dominated by the signal for $N_{p}={\cal O}(10^{4})$ would be needed. This forecast is similar, although more pessimistic, than what is reported in Ref. Cruz:2024svc (see also Tasinato:2023zcg ) due to their extrapolation of the weak-signal limit, which is not expected to hold within the future PTA configurations needed to reach such a level of precision.
@@ V.4 Cosmic Variance
PTA experiments operate in a regime where neither the time average over observation time nor the volume average over the effective detector volume, set by the wavelength of the GWs, allows sampling many realizations of a GWB. Consequently, measurements may be sizeably affected by cosmic variance. This notably affects the measurement of the angular cross-correlation (HD curve), see Refs. Allen:2022dzg ; Bernardo:2022xzl ; Allen:2022ksj ; Allen:2024uqs . In this subsection, we clarify the role of cosmic variance in measuring anisotropies. For simplicity, we will focus on the strong signal limit.
@@ V.5 Hellings-Downs correlation
It is well known that for an anisotropic GWB, the observed angular correlation function scatters around the HD function (see e.g. Taylor:2013esa ). After averaging over many pulsar pairs, the HD function is recovered as the mean value [^9] This implies taking an ensemble average, whereas any actual measurement will be limited by cosmic variance. of the angular correlation function Hotinli:2019tpc , however with an increased variance reflecting the presence of anisotropies Bernardo:2023jhs ; Agarwal:2024hlj ; Allen:2024bnk ; Grimm:2024lfj ; Allen:2024mtn . This implies on the one hand, that the measurement of the HD correlation (averaged over many pulsars) is largely insensitive to the presence of anisotropies. On the other hand, this implies that the HD function is a robust criterion for a GW signal, both for isotropic GWBs, isolated sources, and anisotropic signals arising from multiple point sources Cornish:2013aba ; Cornish:2015ikx ; Taylor:2020zpk ; Becsy:2022pnr ; Allen:2022dzg .
@@ Acknowledgements.
We thank Nihan Pol for the useful discussion and clarifications on the NANOGrav search for anisotropies and Fabrizio Rompineve for collaboration on the initial stages of the project. We thank Bruce Allen for valuable discussions on the HD cosmic variance and in-depth comments on the draft. Moreover, we thank Stanislav Babak, and Carlo Contaldi for the insightful discussions. We thank Ben Farr and Andrea Mitridate for comments on the draft. MP thanks Rutger van Haasteren for the useful discussion on PTA data analysis. MP acknowledges the hospitality of Imperial College London, which provided office space during some parts of this project.
@@ Appendix A Approximations in the derivation of the covariance matrix in the frequency domain
In the following, we assume that there are $N$ observations at $t_{j}=-T_{\text{obs}}/2+jT_{\text{obs}}/N$, $j=0,\dots,N-1$, of all $\Delta t_{I}$ with pulsar index $I$ from 0 to $N_{p}$. Eq. (7) in the main text gives the time delay induced by a GW compared to propagation of the pulsar signal in an Minkowski background. At frequencies $f\gg 1/T_{\text{obs}}$, fitting the timing model to the full dataset achieves this by effectively averaging over many GW oscillations. On the contrary, at low frequencies, $f\sim 1/T_{\text{obs}}$, one is left with a residual sensitivity to the unknown GW phase at $t=0$. To illustrate this, we assume a timing model to be calibrated to $\Delta t(t=0)=0$, leading to
$\Delta t_{I}(t)=\frac{1}{2}\frac{\hat{p}_{I}^{a}\hat{p}_{I}^{b}}{(1+\hat{p}_{I}\cdot\hat{k})}\int_{0}^{t}{\rm d}t^{\prime}\,\left[h_{ab}\left(t^{\prime},\vec{0}\right)-h_{ab}\left(t^{\prime}-D_{I},\hat{p}_{I}D_{I}\right)\right]\,.$ (37)
From this, we can compute the covariance matrix following the sthe onset of the observation (ame steps as in Sec. II.2 of the main text,
$C_{t,h,IJ,ij}$ $=\langle\Delta t_{I}(t_{i})\Delta t_{J}(t_{j})\rangle=\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\left(e^{2\pi ift_{i}}-1\right)\left(e^{-2\pi ift_{j}}-1\right)\,,$ (38)
where $\chi_{IJ}$ is the HD correlation reported in Eq. (15). Compared to Eq. (10) in the main text, we note the extra ‘-1’ terms arising from the ‘miscalibration’ of the pulsar timing model. Upon moving to the frequency domain, and assuming $S_{h}(f)/f^{2}$ to be approximately constant, these terms will contribute only to the $f_{i}=0$ mode (as expected from their role as free integration constant in Eq. (37)) and thus not impact the analysis within the PTA band (see also related discussion in Ref. vanHaasteren:2012hj ). We thus proceed with the expression for the covariance matrix used in the main text,
$C_{t,h,IJ,ij}$ $=\langle\Delta t_{I}(t_{i})\Delta t_{J}(t_{j})\rangle=\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}e^{2\pi if(t_{i}-t_{j})}\,.$
In the case of anisotropies in the GWB, this correlation must be replaced by the overlap reduction function $\Gamma_{IJ}$ given in Eq. (11). Defining the vector of all timing residuals
@@
$=\frac{T_{\text{obs}}^{2}}{N^{2}}\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\;\frac{e^{-i\pi T_{\text{obs}}(f-f_{k})}-e^{i\pi T_{\text{obs}}(f-f_{k})}}{1-e^{i2\pi T_{\text{obs}}(f-f_{k})/N}}\;\frac{e^{i\pi T_{\text{obs}}(f-f_{l})}-e^{-i\pi T_{\text{obs}}(f-f_{l})}}{1-e^{-i2\pi T_{\text{obs}}(f-f_{l})/N}}\,.$ (46)
In PTAs, observation time and pulsar white noise limit the observation band to $f_{\text{min}}\sim 1/T_{\text{obs}}<f<f_{\text{max}}$ with $f_{\text{max}}\sim 10f_{\text{min}}\ll Nf_{\text{min}}$. Consequently, we can approximate $T_{\text{obs}}(f-f_{k})/N\ll 1$ to obtain
$C_{h,IJ,kl}$ $\simeq T_{\text{obs}}^{2}\chi_{IJ}\int_{-\infty}^{\infty}{\rm d}f\frac{S_{h}(f)}{24\pi^{2}f^{2}}\mathrm{sinc}\left[\pi T_{\text{obs}}(f-f_{k})\right]\mathrm{sinc}\left[\pi T_{\text{obs}}(f-f_{l})\right]$ (47)
$\simeq T_{\text{obs}}\,\chi_{IJ}\,\frac{S_{h}(f)}{24\pi^{2}f^{2}}\mathrm{sinc}\left[(f_{k}-f_{l})\pi T_{\text{obs}}\right]$ (48)
$\simeq\chi_{IJ}\frac{S_{h}(f_{k})}{24\pi^{2}f_{k}^{2}}\delta(f_{k}-f_{l})\,.$ (49)
In the second step, we have assumed $S_{h}(f)/f^{2}$ to be approximately constant (otherwise analogous results can be obtained after performing an appropriate whitening procedure). For the third step, recall that $f_{k}$ take integer multiples of $1/T_{\text{obs}}$, which are zeros of the sinc function for any $k\neq l$. Finally, given that Eq. (49) is symmetric in positive and negative frequencies, we can restrict ourselves to positive frequencies only and obtain